# Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding

### v1

@@ 3.1 Data Leakage
Modern LLMs, pre-trained on vast internet data, risk being exposed during training to information about commercial films, such as synopses, reviews, scripts, and transcripts. This issue may result in models recalling answers directly without analyzing the video content, leading to biased benchmarks. In this section, we quantitatively assess the extent of data leakage on datasets. Specifically, we prompt LLMs to answer open-ended questions using only the movie title and compute the accuracy of responses, following [37]. We experiment with 3 movie datasets: MovieQA [56], LVU [64] and our SFD, 5 open-sourced models: Gemma 2B [57], Mistral 7B/8x7B [25], LLaMA-3 8B/70B [58] and 4 commercial models Claude 3 Haiku/Sonnet, and GPT-3.5/4 [42]. Figure 3 reports the results of the data leakage experiment and further details can be found in Appendix G.
@@
We observe that both MovieQA and LVU suffer from data leakage, reaching up to 71.3% and 76.0% accuracies. In contrast, thanks to the low presence of amateur film on the internet, SFD exhibits a maximum accuracy of 36.0%, indicating a low leakage issue. Furthermore, as expected, the extent of data leakage correlates with the knowledge level reflected by MMLU score [14] (see Appendix section G for more experiments), indicating that LLMs with more knowledge exhibit higher levels of memorization, leading to worse leakage issues. For instance, for both MovieQA and LVU, the zero-shot accuracy increases from approximately random level (20%) to more than 70% as the model size rises. Meanwhile, SFD maintains stable and low accuracies ranging between 19.7% and 36.0%, regardless of LLM knowledge variation, further indicating its low data leakage. This experiment reveals that relying solely on existing datasets to evaluate new methods is insufficient. Instead, our SFD offers a more objective and reliable test bed for long-term video understanding.
@@ 3.2 Temporal Window Study
To confirm that our tasks require long-form video understanding, we conduct a temporal window study with benchmarks at three levels: shot-, scene-, and movie-level. A shot, defined as a continuous video clip, typically amounts to around 151 per movie. At the Shot-Level, the model uses data from a single shot, including partial subtitles and visual data, with predictions aggregated by taking the maximum logits. Scene-Level inference aggregates data from approximately 10 shots in a similar manner. Movie-Level inference utilizes the entire film’s data, as detailed in Section 3.3. To further analyze the behaviors of each modality, we ablate at each temporal level three combinations of modalities: Vision, Language, and multimodal Vision-Language. This study is conducted with two top-performing methods from Section 3.3: FrozenBiLM and LLoVi. Further details on the experiment can be found in Appendix E.
@@
Our results, shown in Figure 7, reveal the following trends: (1) Both Language-only and Vision-Language settings show substantial improvements with larger temporal windows. For instance, in the language-only setting, LLoVi’s accuracy increases from 38.5% at the shot-level to 64.2% at the movie-level, reflecting a 25.7% gain. Similarly, FrozenBiLM shows an increasing trend but with a smaller gain in accuracy (+7.3%). The multimodal setting of LLoVi’s accuracy increases from 50.1% to 55.6% at the scene-level but shows only marginal gains at the movie-level. This plateau may be attributed to LLoVi’s naive fusion approach, which combines visual captions with subtitles, limiting further improvements; (2) The Vision-only modality exhibits low and stagnant performance across all levels suggesting that the current handling of visual data is inadequate for this task. Additional results for other methods are presented in Appendix E.
In conclusion, broader temporal windows significantly enhance SFD task performance, especially in language-only settings, underscoring the importance of long-form understanding in our design.
3.3 Baseline Comparison
In this section, we benchmark recent videoQA methods on SFD and evaluate 7 open-source models in three categories: FrozenBiLM [5], mPLUG-Owl2 [74] and Video-LLaVA [33] are VLMs that project visual frames into embedding of an LLM with the help of adapters; MovieChat [54] and TimeChat [47] have more advanced and specific settings, the former introduces a memory mechanism to handle long sequences and the latter proposes a timestamp-aware frame encoder for time-sensitive tasks; LLoVi [79] and LangRepo [26] are text-based methods: they both rely on visual captions on top of an LLM. LangRepo proposes an extra summarization and chain-of-thoughts [62] reasoning prompting.
These models are tested in a zero-shot video question-answering setting, adapted for long-form video understanding by incorporating subtitles and sampling the maximum number of frames possible. In the Multiple-Choice Question answering (MCQ) benchmark, we calculate the accuracy score based on the logits of the predicted options where applicable [5, 33, 74]. For models that generate plain-text answers, we evaluate them using a method similar to the Open-Ended Question answering (OEQ) where we rely on GPT-3.5 to compute the similarity between the predicted and correct answers, as detailed in [37]. All methods are evaluated across three modalities to better assess the contribution of each: Vision-Only (V) with video frames, Language-Only (L) with subtitles, and Vision-Language (VL) combining both. See Appendix D for more details. The results are summarized in Table 3.
User Study. To verify the answerability of our multiple-choice questions and assess the upper limit of SFD, we conducted three user studies accordingly to the aforementioned modality definitions: (1) Vision-Language (VL),—full video with audio and subtitles; (2) Vision-only (V)-muted videos; (3) Language-only (L)-plain text subtitles. For each question, all participants were asked to select the correct answer. Table 3 (last row) reports the results. More details on the user study are available in Appendix F. We observe that when provided with the full multimodal information, participants answer questions with high accuracy (89.8%). As expected, removing modalities lowers accuracy. Specifically, when using only subtitles the performance is 70.9%, whereas the vision-only performance drops to 59.0%.
Method | Venue | % Accuracy | |||||
Multiple-Choice QA | Open-Ended QA | ||||||
V | L | VL | V | L | VL | ||
Random | 20.0 | 20.0 | 20.0 | - | - | - | |
FrozenBiLM [5] | NeurIPS 2021 | 23.4 | 38.2 | 38.6 | - | - | - |
mPLUG-Owl2 [74] | CVPR 2024 | 38.3 | 20.7 | 21.3 | 22.1 | 1.8 | 1.6 |
Video-LLaVA [33] | arXiv 2023 | 34.2 | 21.3 | 24.7 | 19.2 | 6.4 | 8.0 |
LLoVi [79] | arXiv 2023 | 30.8 | 64.2 | 55.6 | 16.2 | 40.3 | 24.7 |
LangRepo [26] | arXiv 2024 | 29.0 | 32.1 | 31.0 | 3.5 | 10.4 | 9.5 |
MovieChat [54] | CVPR 2024 | 8.4 | 6.4 | 8.0 | 14.0 | 15.7 | 11.8 |
TimeChat [47] | CVPR 2024 | 25.5 | 6.4 | 31.8 | 26.4 | 9.4 | 5.9 |
Human | 59.0 | 70.9 | 89.8 | - | - | - |
Multiple-Choice Question answering (MCQ). These results highlight significant insights. Overall, model performance is generally poor, with LLoVi being a notable exception in the language-only setting (64.2%), driven largely by subtitles. Some models, such as FrozenBiLM, perform adequately in the vision-only setting (38.3%), but their performance drops when combining both modalities, indicating integration issues.
Compared to human performance, there is still substantial room for improvement. Accuracy is on par in the language-only setting, with only a 6.7% difference. However, as the best models mainly rely on subtitles, there is a significant gap in the multi-modal setting, where human performance reaches 89.8%. This discrepancy can be explained by the vision-only setting, where the gap reaches 20.7%. These results highlight that while models perform reasonably well with textual data, they struggle to effectively integrate visual information, underscoring the need for advancements in multi-modal fusion techniques to bridge the gap between human and model performance.
Open-Ended Question answering (OEQ). This task proves even more challenging, with lower accuracy across all models, ranging from 3.5% to 40.3%. Subtitles-only models, particularly LLoVi, lead in performance, highlighting the dominance of language processing in understanding and answering complex questions. These findings emphasize the difficulty of the task and the need for improved methods to handle and integrate multi-modal data effectively.
Overall, our results suggest that (1) text (in the form of subtitles) is a stronger cue than visual information for movie question-answering; (2) modern multimodal methods (max accuracy at 60.0%, see Table 3) fall behind human evaluation; and more importantly (3) there is a large room for improvement in the visual aspect, where the gap in performance between modern methods (average accuracy 38.3%) and user study (59.0% accuracy) is still very high.
4 Related Work
VideoQA benchmarks. Questions in such benchmarks can be framed to assess the reasoning, memory, and comprehension skills of both humans and algorithms. Hence, many datasets have been made available through the years, covering visual descriptions [69, 8], temporal action reasoning [65], compositional reasoning [19], social intelligence [76], instructional videos [71], egocentric videos [39, 24, 34], movies [56, 29, 38], among others [23, 67, 23, 32, 75, 63]. However, these datasets feature short videos and typically require reasoning on only a few frames to solve the task at hand, as shown by temporal certificates in [39].
VideoQA methods. To address video question answering, most modern approaches are based on large-scale pre-trained models, by deploying either multimodal contrastive learning [15, 31, 61, 78, 77, 10, 9] or visually-conditioned large language models [30, 11, 68, 74, 2, 70, 73]. For instance, mPLUG-Owl2 [68] enhances LLMs with visual capabilities, while FrozenBiLM [73], a frozen bidirectional language model, uses lightweight adapters and masked language modelling to excel in zero-shot tasks. FrozenBiLM excels in zero-shot video question-answering without using explicit supervision. More recently, Video-LLaVA [33] used a simple projection layer to connect a visual encoder and a large language model, fine-tuned with visual instruction tuning data.
Long-form video understanding. As new long-context models emerge [46], we need more complex benchmarks. Long-form video understanding evaluates the long-term reasoning capabilities of models. Some datasets have been introduced in recent years, mostly for egocentric videos [39] and movies [64, 53, 21, 56]. Instead, LVU [64] involves long movie clips but the subsequent tasks are relatively limited, focusing on irrelevant aspects like ’like’ ratio and view count prediction. [39] proposes multiple-choice questions on 3-minute videos that require on average 100 seconds of video watching to be answered correctly. [53, 21] introduce an audio description task, which is currently tackled at the clip level but could be extended over the whole movie. Finally, closer to our work, [56] and [54] propose a movie question-answering datasets focused on movie plots, settings and characters. However, they do not release any video data because of copyright issues.

### v2

@@ 4 Experiments
In this section, we introduce several experiments performed on the SF20K dataset. Section 4.1 examines data leakage in movie datasets. Section 4.2 assesses human performance on SF20K-Test. Section 4.3 benchmarks recent models across modalities. Section 4.4 reports performance on SF20K subsets. Section 4.5 evaluates temporal window effects on model accuracy. Finally, Section 4.6, analyzes the impact of instruction tuning on VLM performance.
@@ 4.1 Data Leakage
Modern LLMs, pre-trained on vast internet data, risk being exposed during pretraining to information about movies, such as synopses, reviews, scripts, and subtitles. This issue may result in models recalling answers directly without analyzing the video content, leading to biased benchmarks. We call this phenomenon data leakage. In this section, we quantitatively assess the extent of data leakage in movie datasets. For this purpose, we prompt LLMs to answer open-ended questions using only the movie title and compute the response’s correctness following [39]. We experiment with three movie datasets: MovieQA [61], LVU [70], and SF20K-Test; five open-source models: Gemma 2B [62], Mistral 7B/8x7B [28], LLaMA-3 8B/70B [63]; and four commercial models: Claude 3 Haiku/Sonnet [4], GPT-3.5/4 [45]. See Appendix C for more details.
@@
In Figure 3, we observe that both MovieQA and LVU suffer from data leakage, reaching up to 71.3% and 76.0% accuracies. In contrast, thanks to the low presence of amateur films on the internet, SF20K-Test exhibits a maximum accuracy of 36.0%, indicating a low leakage issue. More importantly, the extent of data leakage correlates with the knowledge level reflected by MMLU score [16], indicating that LLMs with more knowledge exhibit higher levels of memorization, leading to worse leakage issues. For instance, for both MovieQA and LVU, the zero-shot accuracy increases from approximately 20% to more than 70% as the LLM size rises. In contrast, SF20K-Test maintains stable and low accuracies ranging between 19.7% and 36.0%, regardless of LLM knowledge variations, further indicating limited knowledge of amateur movies.
| SF20K-Test | MovieQA | |||
| No Anonymization | 31.5 | 71.3 | ||
| Movie Title Removed | 32.8 | +1.3 | 53.9 | -17.4 |
| Full Anonymization | 28.1 | -3.4 | 31.4 | -39.9 |
Anonymization and shuffling. To better understand the impact of movie titles on LLM performance, we perform two additional experiments.
(i) Anonymization. We examine the zero-shot performance of GPT-4 in three settings: (i) providing movie titles, (ii) removing movie titles, and (iii) removing movie titles with additional question anonymization. Question anonymization means replacing character names in the questions with generic placeholders, as this may help LLMs to recall specific movies (e.g., Harry Potter, James Bond). It enables us to measure relative accuracies, i.e. the performance improvement of an LLM if it can guess the movie. For MovieQA, removing movie titles causes a significant drop in performance (-17.4%), with further reduction after full anonymization (-39.9%), as shown in Table 2. In sharp contrast, the performance on the SF20K-Test remains overall stable. This indicates that the ability to identify the movie is the main reason why LLMs perform so well in commercial movies in our data leakage experiment, while this does not apply to amateur films.
(ii) Shuffling. We examine the data leakage performance of Llama-3-70B in a scenario where movie titles are replaced with random titles from the same dataset instead of providing the correct ones. As shown in Table 3, this substitution causes a substantial drop in accuracy for MovieQA, even falling below the fully anonymized setting. This suggests that LLMs rely heavily on recognizing commercial movie titles for their predictions. On the contrary, for SF20K-Test, replacing correct titles with random ones does not notably affect performance, likely because the LLM does not recognize these amateur films in either case.
These experiments show that (i) evaluating new methods on commercial movies can be misleading, and (ii) SF20K-Test provides a more reliable benchmark for long-term video understanding.
| SF20K-Test | MovieQA | |||
| Correct movie title | 24.1 | 51.9 | ||
| Random movie title | 19.2 | -4.9 | 10.9 | -41.0 |
4.2 User study
To verify the answerability of our multiple-choice questions and assess the upper limit of SF20K-Test, we conduct three user studies according to the following modality definitions: (1) Vision-Language (VL), i.e. full video with audio and subtitles; (2) Vision-only (V), i.e. muted videos; (3) Language-only (L), i.e. plain text subtitles. For each question, all participants are asked to select the correct answer (see Appendix D for more details). Table 4 (last row) reports the results. We observe that when provided with the full multimodal information, participants answer questions with high accuracy (89.8%). As expected, removing modalities lowers accuracy. Specifically, when using only subtitles, the performance drops to 70.9%, whereas the vision-only performance decreases to 59.0%.
4.3 Baselines
In this section, we evaluate recent open-source and commercial vision-language models on SF20K-Test. In particular, we benchmark eleven open-source models: FrozenBiLM [6], mPLUG-Owl2 [80], TimeChat [52], Video-Llava [36], MovieChat [59], Llava-OneVision [32], Long-Llava [67], LongVA [84], Llava-Next-Video [85], Llama-3.2-Vision [43], and Pixtral [3]; and two commercial models: LloVi [83] (based on GPT-3.5 [45]), and GPT-4o-mini [46]. See Appendix E for more details on prompts and metrics.
The models are tested in a zero-shot video question-answering setting, adapted for long-form video understanding by incorporating subtitles and sampling the maximum number of frames possible to fit in a single A100 GPU. For multiple-choice questions, we compute the accuracy score by extracting the option letter choice from the model’s response. For open-ended questions, we rely on GPT-3.5 to compute the similarity between the predicted and correct answers, as detailed in [39]. All methods are evaluated across three modalities to better assess the contribution of each: Vision-Only (V) with video frames, Language-Only (L) with subtitles, and Vision-Language (VL) combining both. The results are summarized in Table 4.
| Method | Multiple-Choice QA | Open-Ended QA | ||||
| V | L | VL | V | L | VL | |
| Random | 20.0 | 20.0 | 20.0 | - | - | - |
| Open-source models | ||||||
| FrozenBiLM [79] | 23.4 | 38.2 | 38.6 | - | - | - |
| mPLUG-Owl2 [80] | 38.3 | 20.7 | 21.3 | 22.1 | 1.8 | 1.6 |
| TimeChat [52] | 25.5 | 6.4 | 31.8 | 26.4 | 9.4 | 5.9 |
| Video-Llava-7B [35] | 32.4 | 21.3 | 45.7 | 8.4 | 10.6 | 8.2 |
| MovieChat [59] | 8.4 | 6.4 | 8.0 | 14.0 | 15.7 | 11.8 |
| Llava-OneVision-0.5B [32] | 29.2 | 35.4 | 37.8 | 14.3 | 20.2 | 22.6 |
| Long-Llava-7B [67] | 52.1 | 70.4 | 72.6 | 17.4 | 36.7 | 33.5 |
| LongVA-7B [84] | 58.5 | 73.8 | 70.0 | 26.3 | 51.7 | 36.6 |
| Llava-OneVision-7B [32] | 63.2 | 74.7 | 78.2 | 25.1 | 35.3 | 37.7 |
| Llava-Next-Video-7B [85] | 32.8 | 52.3 | 64.2 | 27.4 | 42.2 | 42.7 |
| Llama-3.2-Vision-11B [43] | 57.9 | 72.3 | 75.0 | 24.4 | 45.2 | 47.8 |
| Pixtral-12B [3] | 38.5 | - | 74.4 | 29.5 | 49.4 | 49.8 |
| Commercial models | ||||||
| LloVi [83] (GPT-3.5-based) | 30.8 | 64.2 | 55.6 | 16.2 | 40.3 | 24.7 |
| GPT-4o-mini [46] | - | - | 79.4* | - | - | 60.0* |
| Human | 59.0 | 70.9 | 89.8 | - | - | - |
∗ indicates that inference is done on only 10% of the test set to limit costs.
Multiple-Choice Question Answering (MCQA). Our results reveal that several models exhibit strong performance, with GPT-4o-mini achieving the highest accuracy of 79.4%. Notably, recent open-source models are closing the gap with commercial ones; for example, Llava-OneVision-7B lags behind by just 1.2 percentage points. However, there remains significant room for improvement, as human performance reaches 89.8%.
Open-Ended Question Answering (OEQA). This task is more challenging, as evidenced by significantly lower accuracy across all models, ranging from 1.6% to 60.0%. GPT-4o-mini again leads in performance. On the other hand, open-source models face a larger performance gap. For instance, Pixtral-12B achieves a maximum accuracy of only 49.8%. Appendix F provides an analysis of failure cases (see Table 12).
Modality Ablations. For both tasks, recent models demonstrate improved modality integration, as multimodal performance consistently surpasses unimodal performance. However, language remains the dominant modality, as evidenced by higher accuracy in language-only settings compared to vision-only settings, and a relatively small performance gap between language-only and multimodal settings (+19.9% and +0.4%, respectively, for Pixtral-12B on open-ended questions).
Overall, our results indicate that (1) recent models perform reasonably well on multiple-choice questions, though there remains substantial room for improvement compared to human performance; (2) open-ended question answering poses a much greater challenge, with all models struggling, particularly open-source models who lag behind their commercial counterparts; and (3) performance relies heavily on subtitles, suggesting that better integration of the visual modality could improve accuracy.
| Model | SF20K-Test | Silent | SF20K-Test-Expert | ||||
| V | L | VL | V | V | L | VL | |
| Open-source models | |||||||
| Video-LLaVA-7B [35] | 8.4 | 10.6 | 8.2 | 7.2 | 9.0 | 5.0 | 3.5 |
| Llava-Next-Video-7B [85] | 27.4 | 42.2 | 42.7 | - | 14.2 | 10.3 | - |
| Llava-OneVision-0.5B [32] | 14.3 | 20.2 | 22.6 | 17.6 | 12.5 | 12.0 | 19.5 |
| Llava-OneVision-7B [32] | 25.1 | 35.3 | 37.7 | 29.0 | 21.5 | 20.8 | 29.0 |
| Long-Llava-7B [84] | 17.4 | 36.7 | 33.5 | 20.9 | 17.2 | 22.0 | 29.5 |
| Llama-3.2-Vision-11B [43] | 24.4 | 45.2 | 47.8 | 29.0 | 17.2 | 29.8 | 31.0 |
| Pixtral-12B [3] | 29.5 | 49.4 | 49.8 | 17.0 | 17.0 | 30.9 | 33.0 |
| Commercial models | |||||||
| GPT-4o-mini [2] | - | - | 60.0* | - | - | - | - |
∗ indicates that inference is done on only 10% of the test set to limit costs.
4.4 Performance on SF20K-Test subsets
In this section, we evaluate recent VLMs on two subsets: SF20K-Test-Silent (silent movies, without dialogs) and SF20K-Test-Expert (manual questions created by expert annotators). Both subsets are evaluated in an OEQA setting. Table 5 reports the results. For silent movies, the modality ablation is not performed since there are no transcripts available.
SF20K-Test-Silent. Interestingly, all models present a significant performance drop in the silent subset compared to the normal subset, with accuracy not exceeding 30%. Note that these questions are similar in type and difficulty to those in the full SF20K-Test, where models reach 49.8% accuracy. For example, Llama-3.2-Vision-11B performance declines by 18.8%, and other models show similar trends. Moreover, performance differences between models are reduced in this subset, with both Llava-OneVision-7B and Llama-3.2-Vision-11B now ranking as top-performing models. This consistent drop in accuracy suggests that models are not effectively leveraging visual information alone.
SF20K-Test-Expert. On challenging questions, all models experience a performance drop, with both Pixtral-12B and Llama-3.2-Vision-11B showing a 16.8% decline. Moreover, we observe that the Language-only (L) and Vision-Language (VL) performances are similar, suggesting that subtitles (L) continue to be the primary contributor to model performance, which indicates the models’ limited reliance on visual information.
| Fine-tuned | Multiple-Choice QA | Open-Ended QA | ||||||||||
| V | L | VL | V | L | VL | |||||||
| ✗ | 29.2 | 35.4 | 37.8 | 14.3 | 20.2 | 22.6 | ||||||
| ✓ | 45.7 | +16.5 | 61.2 | +25.8 | 63.0 | +25.2 | 17.8 | +3.5 | 22.5 | +2.3 | 26.6 | +4.0 |
@@ 4.5 Temporal Window Study
To confirm that our tasks require long-form video understanding, we conduct a temporal window study by performing inference at three levels: shot-, scene-, and movie-level. Shots, defined as continuous video clips, last 5.52 seconds on average. At the Shot-Level, the model uses data from one shot, including partial subtitles and visual data, with predictions aggregated by taking the maximum logits over all shots. Scene-Level inference aggregates data from approx. 10 shots, in a similar manner. Movie-Level inference utilizes the entire film’s data. To further analyze the behavior of each modality, we ablate at each temporal level three combinations of modalities: Vision, Language, and Vision-Language. This study is conducted with two methods: FrozenBiLM and LLoVi, for the MCQA task. More details in Appendix G.
@@
Our results, shown in Figure 6, reveal the following trends: (1) Both Language-only and Vision-Language settings show substantial improvements with larger temporal windows. For instance, in the language-only setting, LLoVi’s accuracy increases from 38.5% at the shot-level to 64.2% at the movie-level, reflecting a 25.7% gain. Similarly, FrozenBiLM shows an increasing trend but with a smaller gain in accuracy (+7.3%). The multimodal setting of LLoVi’s accuracy increases from 50.1% to 55.6% at the scene-level but shows only marginal gains at the movie-level. This plateau may be attributed to LLoVi’s naive fusion approach, which combines visual captions with subtitles, limiting further improvements; (2) The Vision-only modality exhibits low and stagnant performance across all levels, suggesting that the current handling of visual data is inadequate for this task.
In conclusion, larger temporal windows substantially enhance the performance on SF20K-Test, especially in language-only setting. This underscores the long-term nature of our proposed story-level video tasks.
4.6 Instruction tuning
In this experiment, we fine-tune Llava-OneVision-0.5B [32] on SF20K-Train (Section 3.2), using LoRA for efficiency. The input includes video frames, subtitles, and the question. We use next-token prediction to generate responses. In the multiple-choice case, we add multiple-choice options to the query and predict the correct option index in addition to the answer. Fine-tuning is performed over 191,007 samples from 19,676 movies for 5 epochs, with a learning rate of 1.0e-4, 50 warmup steps, batch size 8, gradient norm 1.0, and LoRA parameters (r=8, alpha=8, dropout=0.1). Samples are randomly presented as multiple-choice or open-ended questions.
Instruction tuning performance. Table 6 reports the results, showing that instruction tuning consistently enhances performance across all modalities and tasks compared to the pre-trained model. For multiple-choice QA, performance improves significantly by 25.2%, reducing the gap with its 7B counterparts. In the open-ended QA setting, while the improvement is more modest (+4.0%), it remains meaningful. Notably, the vision-only modality also benefits from instruction tuning, despite the instruction data being generated from subtitles.
Effect of the training size. Figure 9 shows the performance of Llava-OneVision-0.5B [32] when instruction-tuned for multiple-choice question answering on SF20K-Train, with varying numbers of training videos. We subsample the full dataset at incremental values. Results indicate a logarithmic improvement in accuracy as the training set size increases, rising from 46.5% to 63.0%. Results are displayed in log scale. These findings highlight the scalability benefits of instruction tuning on SF20K-Train.